OpenStudy (anonymous):

Solve. 15 – z = 6 A. –9 B. 21 C. –21 D. 9

OpenStudy (tkhunny):

Have you considered adding z and subtracting 6?

OpenStudy (anonymous):

is it 21

OpenStudy (tkhunny):

Demonstrate your work and let's see where you wandered off. Show what you get after adding z to both sides.

OpenStudy (anonymous):

i said 15+6

OpenStudy (tkhunny):

No, that's the magic method where you forget all the rules. Just add z to both sides and show your result.

OpenStudy (anonymous):

idk

OpenStudy (tkhunny):

15 - z + z = 15 + 0 = 15 You do the right side.

OpenStudy (anonymous):

9

OpenStudy (tkhunny):

{sigh} - Part of what you should be learning is thinking and styling. Trowing down possible answers is not very helpful. You have to develop a style that is consistent and that actually makes your life easier - rather than a style that leaves you guessing half the time. 15 – z = 6 Add z 15 – z + z = 6 + z Simplify 15 = 6 + z Subtract 6 15 - 6 = 6 + z - 6 Simplify and rearrange 9 = 6 - 6 + z Simplify 9 = z Straight forward. Easily followed. No question what happened or why.